OpenStudy (anonymous):

solve the equation. √2x+15 = √x+10

OpenStudy (anonymous):

the root for 2x , x only or for (2x+15)and (x+10)

OpenStudy (imstuck):

You have to be a bit more specific with what exactly is under that square root sign. Ok?

OpenStudy (anonymous):

the root is for 2x++15 and the root is for x+10

OpenStudy (anonymous):

sorry i dont know how to make it cover them both

OpenStudy (imstuck):

so the whole thing (2x + 15) is under a square root? and then (x + 10) is under a square root too?

OpenStudy (anonymous):

\[\sqrt{2x+15}=\sqrt{x+12} ~~ or ~~ \sqrt{2x}+15=\sqrt{x}+10\]

OpenStudy (anonymous):